x2+1=0

y=x2+1

find the solution of the given system of equations.


1
Expert's answer
2020-12-31T15:17:35-0500

"\pmb{x^2+1=0}\;\;\; ..(1)\\
\pmb{y=x^2+1}\;\;\;..(2)"


Now from equation "(1),"

"x^2+1=0\\
\Rightarrow (x)^2-(-1)=0\\
\Rightarrow (x)^2-(i)^2=0\\
\pmb[where \;\;i=\sqrt{-1}\;\pmb]\\
\Rightarrow (x-i)(x+i)=0\\
\Rightarrow x=i\;\;\;or, \;\;\;x=-i"


Now putting these "x"-values in equation "(2)" ,

"y=0" if "x=i"

and, "y=0" if "x=-i"


"\therefore" From above discussion, we can decide that,

Solution set of "(x,y)=\{(i,0),(-i,0)\}" [Ans.]


i.e. "(x,y)= (i,0) \;\;\;or, \;\;\;(-i,0)"